# Lack of collisional hydrodynamics in a harmonically trapped one-dimensional Bose gas

### v1

@@ Acknowledgements.
The authors thank H. Hu and X.-J. Liu for kindly sharing the CHDVA data from Ref. [30] shown in Fig. 2 for comparison purposes. C.M. is grateful for the computational support of I. Mortimer and the Core Computing Facility “Getafix” of the School of Mathematics and Physics at the University of Queensland. K.V.K. acknowledges support by the Australian Research Council Discovery Project Grant No. DP240101033. G.D.R. received funding from the Universitat Politècnica de Catalunya (ALECTORS2024, R-02673) and the grant IJC2020-043542-I funded by MCIN/AEI/10.13039/501100011033 and by "European Union NextGenerationEU/PRTR". G.D.R. also acknowledges support by the Spanish Ministerio de Ciencia e Innovación (MCIN/AEI/10.13039/501100011033, Grant PID2023-147469NB-C21), and by the Generalitat de Catalunya (Grant 2021 SGR 01411).
@@
Collective oscillations are generated by a perturbation $H_{\rm pert}=-\lambda F(x)\theta(t)$ to the Hamiltonian (1) with a small strength parameter $\lambda$. Here, $F(x)=\sum_{i=1}^{N}f\left(x_{i}\right)$ is the static excitation operator and $\theta(t)$ is the Heaviside function of time $tH_{\rm pert}$ is implemented as a sudden quench of the external trapping potential $V(x)\rightarrow V(x)-\lambda f(x)$ at $t=0$ leading to a density perturbation $n(x,t)$.
@@
Figure 4:  (1): Example of the skewness, Eq. (5), vs. the dimensionless time $t\omega_{x}$, calculated for dataset (c) (see Table 2) at $\gamma_{0}\!=\!5.035$, corresponding to N=1412N\!=\!1412 in Fig. 3(2): Fourier transform of the skewness in (1), $\rm{FT}\left[\rm{Skew}\right]$, Eq. (6), vs. frequency $\omega/\omega_{x}$. The yellow and blue markers at the two peaks correspond to the DC oscillation frequencies $\omega_{1}$ and $\omega_{2}$.

### v2

@@ Acknowledgements.
The authors thank H. Hu and X.-J. Liu for kindly sharing the CHD variational ansatz data from Ref. [29] shown in Fig. 2 for comparison purposes. C.M. is grateful for the computational support of I. Mortimer and the Core Computing Facility “Getafix” of the School of Mathematics and Physics at the University of Queensland. K.V.K. acknowledges support by the Australian Research Council Discovery Project Grant No. DP240101033. G.D.R. received funding from the Universitat Politècnica de Catalunya (ALECTORS2024, R-02673) and the grant IJC2020-043542-I funded by MCIN/AEI/10.13039/501100011033 and by "European Union NextGenerationEU/PRTR". G.D.R. also acknowledges support by the Spanish Ministerio de Ciencia, Innovación y Universidades (grant PID2023-147469NB-C21, financed by MICIU/AEI/10.13039/501100011033 and FEDER-EU).
@@
Collective oscillations are generated by a perturbation $H_{\rm pert}=-\lambda F(x)\theta(t)$ to the Hamiltonian (1) with a small strength parameter $\lambda$. Here, $F(x)=\sum_{i=1}^{N}f\left(x_{i}\right)$ is the static excitation operator and $\theta(t)$ is the Heaviside function of time $t$. $H_{\rm pert}$ is implemented as a sudden quench of the external trapping potential $V(x)\rightarrow V(x)-\lambda f(x)$ at $t=0$ leading to a density perturbation $n(x,t)$.
@@
Figure 4: (1): Example of the skewness, Eq. (5), vs. the dimensionless time $t\omega_{x}$, calculated for dataset (c) (see Table 2) at $\gamma_{0}\!=\!5.035$, corresponding to N=1412N\!=\!1412 in Fig. 3. (2): Fourier transform of the skewness in (1), $\rm{FT}\left[\rm{Skew}\right]$, Eq. (6), vs. frequency $\omega/\omega_{x}$. The yellow and blue markers at the two peaks correspond to the DC oscillation frequencies $\omega_{1}$ and $\omega_{2}$.